Trade-Offs and Trade Work It Out: Question 3 of 4 Assume the economy of Atlantis only produces fish (point F). Calculate the opportunity cost of producing potatoes in 200 pound increments. Quantity of potatoes (pounds) 1,000 800 600 400 Quantity of fish (pounds) Maximum annual output options 300 500 600 650 675 200Opportunity cost of increasing production from 0 pounds to 200 pounds of potatoes: pounds of fish. Opportunity cost of increasing production from 200 pounds to 400 pounds of potatoes: pounds of fish. Opportunity cost of increasing production from 400 pounds to 600 pounds of potatoes: pounds of fish. Opportunity cost of increasing production from 600 pounds to 800 pounds of potatoes: pounds of fish. Opportunity cost of increasing production from 800 pounds to 1,000 pounds of potatoes: pounds of fish.
Answer #1

Opportunity cost is the value of next best alternative foregone.

Opportunity cost of increasing production of potatoes is the amount of fish which is sacrificed for that additional production of potatoes.

1) Opportunity cost = 675 - 650 = 25 pounds of fish

2) Opportunity cost = 650 - 600 = 50 pounds of fish

3) Opportunity cost = 600 - 500 = 100 pounds of fish

4) Opportunity cost = 500 - 300 = 200 pounds of fish

5) Opportunity cost = 300 - 0 = 300 pounds of fish